Returns. What are the returns on the following investments? Original Cost Selling Price of Investment of Investment Distributions Percent Return $100 $23 $160 $10 $1,100 523,000 S0 $2 $80 S0 Bond Car $1,040 $53,000 Original Cost $100 S23 $1,040 Selling PriceDistributions of Percent Retum CD Stock Bond $160 $10 1,100 S0 S2 $80 60 % (Round to two decimal places.) -4783 % (Round to two decimal places.) I 500-1 % (Round to two decimal places )
Answer #1

Return is computed as (Capital gain + distributions)/ Original investment

CD= (160-100+0)/100 = 60%

Stock= (10-23+2)/ 23 = -47.83%

Bond= (1100-1040+80)/ 1040 = 13.46%

Car= (23000-53000)/ 53000 = -566.04%